# Reexamining Paradigms of End-to-End Data Movement

### v4

@@ 3.2.2. The TCP CCA Paradigm
The empirical evidence presented previously established the irrelevance of packet loss on well-engineered network backbones. Next, we examine the related paradigm that the choice of TCP CCA is critical for performance. The prevalence of this view is evident from the abstract of Google Networking Research's publication "BBR: Congestion-Based Congestion Control" (Cardwell et al., 2016), which introduced the BBR algorithm (Bottleneck Bandwidth and Round-trip Propagation Time). The abstract states: “Physics and climate researchers need to exchange petabytes of data with global collaborators but find their carefully engineered multi-Gbps infrastructure often delivers at only a few Mbps over intercontinental distances.”—a claim that implicitly attributes poor performance to TCP behavior rather than end-system or architectural factors.
It is a telling contrast that during the very period this paradigm was being articulated, a production reality was demonstrated. ESnet publicly reported our 1 PB transfer, which achieved 76.63 Gbps average utilization (Energy Sciences Network (2018), ESnet). This result was accomplished not with a novel CCA, but with the default CUBIC in the standard CentOS 7.5 distribution, which did not yet include BBR. This demonstrates that the fundamental bottleneck was not the CCA, but the lack of a holistically co-designed system.
Before proceeding, we establish the context regarding high-speed TCP CCAs by citing relevant Internet Engineering Task Force (IETF) Request for Comments (RFCs) and referring to prior work (Energy Sciences Network (2018), ESnet). Starting in 2018, the suitability of CUBIC for Fast Long-Distance Networks was already documented by RFC8312 (Rhee et al., 2018). Coincidentally, earlier work (Energy Sciences Network (2018), ESnet) described a production trial conducted on a 5000-mile, 100 Gbps loop, achieving over 96% bandwidth utilization—practically, line rate. The systems used in this trial ran CentOS 7.5-1804, which utilized CUBIC as its default CCA. We note that RFC9438 (Xu et al., 2023), titled “CUBIC for Fast and Long-Distance Networks”, has since obsoleted RFC8312. More recently, RHEL 9.6 and its free rebuilds now include BBR as the default CCA, as demonstrated by the simple command output presented below (command line output wrapped to fit in a single column):
@@
  • Hyperscale Data Set: A synthetic test data set containing files of a uniform size, where the total number of files is 220 (1,048,576) or the aggregate size is 1 TiB, or both.
  • Data transfer sweep:
@@ 3.2.3. Practical Implications and Real-World Consequences
This dichotomy recalls the hierarchy of solutions presented in Figs. 2-3, whose importance can be illustrated by a simple metaphor: providing water. Watering plants on a balcony requires a watering can. Watering a garden demands a hose. However, supplying a metropolitan area from a reservoir necessitates a massively engineered system, e.g., the Hetch Hetchy Aqueduct (Bay Area Water Supply and Conservation Agency (2025), BAWSCA).
@@ 3.3. Dedicated Private Lines Are Essential for High-Speed Testing
From 2015-2019, the zx R&D was carried out on a testbed at SLAC National Accelerator Laboratory, using a 5,000-mile, 100 Gbps loop provisioned by ESnet. After 2019, collaboration with ESnet continued (Kissel and Fang, 2020c). Nevertheless, accessing their 100G Software Defined Network (SDN) Testbed required a formal application process (Energy Sciences Network (2025f), ESnet), hindering agile development. This need for a fully controlled, on-demand test environment led us to evaluate commercial WAN simulators, which we found to be prohibitively expensive and functionally limited.
@@ 3.5. Powerful CPUs are essential for high transfer rates
This skepticism was grounded in years of deployment experience, which revealed that zx consistently achieves high performance without requiring a high core count; typically, 12-24 cores are sufficient. While resources like ESnet's fasterdata recommend high clock rates for good encryption performance (Energy Sciences Network (2025c), ESnet), our observations suggested an alternative path. We hypothesized that fewer cores with CPU-built-in encryption acceleration and moderate clock rates would reduce context switches (improving software efficiency) and lower energy consumption, thereby reducing the TCO. This led to the selection of the Intel Xeon 5418N, a mid-range model (Intel, 2025a).
@@ 3.6.4. Empirical Evidence of the Cloud Data Movement Limitations
The performance penalties imposed by standard cloud data paths are measurable. Using the GoToCloud platform on AWS (Moriya et al., 2024), a comparative analysis conducted by High Energy Accelerator Research Organization (KEK) (High Energy Accelerator Research Organization (2025), KEK) vividly illustrates this point. As summarized in Fig. 11, the transfer of a 1.2 TiB Cryo-electron microscopy (Cryo-EM) dataset was measured with zx vs. aws-cli
@@ 5.1. Validation Platforms
This testing is part of a decade-long, sustained effort to validate the co-design principle across diverse production environments, as documented in prior studies (Thayer and Fang, 2019; Energy Sciences Network (2017), ESnet; Energy Sciences Network (2018), ESnet; ICM, 2019; NSCC, 2019; Fang et al., 2016; Fang, 2019).
@@
  • Two mini appliances deployed at two distinct sites in Taiwan, connected via the Taiwan Advanced Research and Education Network (TWAREN) (TaiWan Advanced Research and Education Network (2025), TWAREN).
@@ 6. Conclusions
The scalability of this architectural approach was demonstrated as early as 2018 in the petabyte-transfer setup (Energy Sciences Network (2017), ESnet) and (Energy Sciences Network (2018), ESnet) shown in Fig. 12 (reproduced from (Kissel and Fang, 2020c) Appendix 6.1). This scale-out design used multiple older servers with zx-aggregated 10 Gbps interfaces, proving that the invariant software architecture could distribute workload across nodes without cluster management overhead. This consistency ensures that the logical appliance design—whether implemented as a 10 Gbps mini-appliance or a 400 Gbps core node—remains uniform across scales. For example, modern 2U2N servers, e.g., the SYS- [^2] BT-DNR (Supermicro, 2025) (Fig. 13), enable the same scale-out efficiency within a single chassis. It extends our testbed-based methodology to 400 Gbps and beyond, using the same architectural principles that successfully moved petabytes in 2018.

### v5

@@ 3.2.2. The TCP CCA Paradigm
The empirical evidence presented previously established the irrelevance of packet loss on well-engineered network backbones. Next, we examine the related paradigm that the choice of TCP CCA is critical for performance. The prevalence of this view is evident from the abstract of Google Networking Research's publication "BBR: Congestion-Based Congestion Control" (Cardwell and others, 2016), which introduced the BBR algorithm (Bottleneck Bandwidth and Round-trip Propagation Time). The abstract states: “Physics and climate researchers need to exchange petabytes of data with global collaborators but find their carefully engineered multi-Gbps infrastructure often delivers at only a few Mbps over intercontinental distances.”—a claim that implicitly attributes poor performance to TCP behavior rather than end-system or architectural factors.
It is a telling contrast that during the very period this paradigm was being articulated, a production reality was demonstrated. ESnet publicly reported our 1 PB transfer, which achieved 76.63 Gbps average utilization (Energy Sciences Network (ESnet), 2018). This result was accomplished not with a novel CCA, but with the default CUBIC in the standard CentOS 7.5 distribution, which did not yet include BBR. This demonstrates that the fundamental bottleneck was not the CCA, but the lack of a holistically co-designed system.
Before proceeding, we establish the context regarding high-speed TCP CCAs by citing relevant Internet Engineering Task Force (IETF) Request for Comments (RFCs) and referring to prior work (Energy Sciences Network (ESnet), 2018). Starting in 2018, the suitability of CUBIC for Fast Long-Distance Networks was already documented by RFC8312 (Rhee and others, 2018). Coincidentally, earlier work (Energy Sciences Network (ESnet), 2018) described a production trial conducted on a 5000-mile, 100 Gbps loop, achieving over 96% bandwidth utilization—practically, line rate. The systems used in this trial ran CentOS 7.5-1804, which utilized CUBIC as its default CCA. We note that RFC9438 (Xu and others, 2023), titled “CUBIC for Fast and Long-Distance Networks”, has since obsoleted RFC8312. More recently, RHEL 9.6 and its free rebuilds now include BBR as the default CCA, as demonstrated by the simple command output presented below (command line output wrapped to fit in a single column):
@@
  • Hyperscale Data Set. A synthetic test data set containing files of a uniform size, where the total number of files is $2^{20}$ (1,048,576) or the aggregate size is 1 TiB, or both.
  • Data Transfer Sweep.
@@ 3.2.3. Practical Implications and Real-World Consequences
This dichotomy recalls the hierarchy of solutions presented in Figs. 2-3, whose importance can be illustrated by a simple metaphor: providing water. Watering plants on a balcony requires a watering can. Watering a garden demands a hose. However, supplying a metropolitan area from a reservoir necessitates a massively engineered system, e.g., the Hetch Hetchy Aqueduct (Bay Area Water Supply and Conservation Agency (BAWSCA), 2025).
@@ 3.3. Dedicated Private Lines Are Essential for High-Speed Testing
From 2015-2019, the zx R&D was carried out on a testbed at SLAC National Accelerator Laboratory, using a 5,000-mile, 100 Gbps loop provisioned by ESnet. After 2019, collaboration with ESnet continued (Kissel and Fang, 2020c). Nevertheless, accessing their 100G Software Defined Network (SDN) Testbed required a formal application process (Energy Sciences Network (ESnet), 2025g), hindering agile development. This need for a fully controlled, on-demand test environment led us to evaluate commercial WAN simulators, which we found to be prohibitively expensive and functionally limited.
@@ 3.4. Increasing network bandwidth increases transfer rates
Closely related and often overlooked is the dataset’s file size distribution. Workloads dominated by large numbers of small files incur substantial per-file overheads—including metadata operations, open/close costs, and short-lived transfers—that disrupt effective pipelining and prevent full utilization of available bandwidth. This effect is evident in Figs. 56 and 89, where increased parallelism or bandwidth does not translate into proportional throughput gains. In such cases, the system repeatedly fails to reach steady-state data flow, so adding bandwidth alone yields little benefit. Note also that workloads composed of only a few extremely large files present a different challenge. Because most data movers use the file as the fundamental I/O scheduling unit, the achievable concurrency is effectively bounded by the number of files. When datasets contain only a handful of multi-hundred-GiB or TiB-scale files, sufficient parallelism cannot be attained to saturate high-bandwidth links. Even aggressive tuning cannot compensate for the limited number of independent I/O operations. This effect has been observed in operational DTN environments (Kissel and Fang, 2020c) where transfer rates of the baseline data mover decline as file sizes grow and file counts shrink. In such cases, the bottleneck is not per-file overhead but concurrency starvation.
@@ 3.5. Powerful CPUs are essential for high transfer rates
This skepticism was grounded in years of deployment experience, which revealed that zx consistently achieves high performance without requiring a high core count; typically, 12-24 cores are sufficient. While resources like ESnet's fasterdata recommend high clock rates for good encryption performance (Energy Sciences Network (ESnet), 2025c), our observations suggested an alternative path. We hypothesized that fewer cores with CPU-built-in encryption acceleration and moderate clock rates would reduce context switches (improving software efficiency) and lower energy consumption, thereby reducing the TCO. This led to the selection of the Intel Xeon 5418N, a mid-range model (Intel, 2025a).
@@ 3.6.4. Empirical Evidence of the Cloud Data Movement Limitations
The performance penalties imposed by standard cloud data paths are measurable. Using the GoToCloud platform on AWS (Moriya and others, 2024), a comparative analysis conducted by High Energy Accelerator Research Organization (KEK) (High Energy Accelerator Research Organization (KEK), 2025) vividly illustrates this point. As summarized in Fig. 11, the transfer of a 1.2 TiB Cryo-electron microscopy (Cryo-EM) dataset was measured with zx vs. aws-cli
@@ 5.1. Validation Platforms
This testing is part of a decade-long, sustained effort to validate the co-design principle across diverse production environments, as documented in prior studies (Thayer and Fang, 2019; Energy Sciences Network (ESnet), 2017, 2018; ICM, 2019; NSCC, 2019; Fang et al., 2016; Fang, 2019).
@@
  • Two mini appliances deployed at two distinct sites in Taiwan, connected via the Taiwan Advanced Research and Education Network (TWAREN) (TaiWan Advanced Research and Education Network (TWAREN), 2025).
@@ 6. Conclusions
The scalability of this architectural approach was demonstrated as early as 2018 in the petabyte-transfer setup (Energy Sciences Network (ESnet), 2017) and (Energy Sciences Network (ESnet), 2018) shown in Fig. 12 (reproduced from (Kissel and Fang, 2020c) Appendix 6.1). This scale-out design used multiple older servers with zx-aggregated 10 Gbps interfaces, proving that the invariant software architecture could distribute workload across nodes without cluster management overhead. This consistency ensures that the logical appliance design—whether implemented as a 10 Gbps mini-appliance or a 400 Gbps core node—remains uniform across scales. For example, modern 2U2N servers, e.g., the SYS- [^2] BT-DNR (Supermicro, 2025) (Fig. 13), enable the same scale-out efficiency within a single chassis. It extends our testbed-based methodology to 400 Gbps and beyond, using the same architectural principles that successfully moved petabytes in 2018.